# Multi-Agent Taskforce Collaboration: Self-Correction of Compounding Errors in Long-Form Literature Review Generation

### v2

@@ 3.3 Feedback Taskforce
The feedback taskforce is designed to execute tasks assigned by the manager, taking into account the risk of deviations at each step arising from unpredictable agent-environment interactions. To mitigate compounding errors, we propose an experience-based strategy.
@@
In detail, for each agent $A_{W}$, the manager agent $A_{M}$ initiates the feedback taskforce. The agent first generates the primary results, then $A_{M}$ leverages historical experience to review these results, after which $A_{W}$ refines its outputs based on the feedback provided. At the initial step, $A_{W}$ produces the first output solely based on the context:
@@
where $Y^{(0)}$ is the initial output, and $C$ is the context fetched by the manager agent.
@@
To give useful feedback, we propose an experience-aware mechanism. We record feedback and revision history in a database and select useful feedback for demonstration to provide feedback on the worker’s output at each iteration $x$.
$r^{*}=\arg\max_{r\in\mathcal{R}_{\min}}\text{usage}(r)$ (12)
where useful feedback is defined as the feedback that has led to the fewest further revisions, where $\mathcal{R}_{\min}=\{r\in\mathcal{R}:\text{revision}(r)=r_{\min}\}$. Here, $\text{revision}(r)$ denotes the number of revisions after using feedback $r$. In $\mathcal{R}_{\min}$, we select the most useful one, i.e., with the highest $\text{usage}(r)$. Here, $\text{usage}(r)$ denotes the number of times the feedback $r$ is selected for demonstration.
Referring to $r^{*}$, the manager agent then provides feedback on the worker’s output:
$R^{(x)}=A_{M}(Y^{(x)},r^{*})$ (13)
where $R^{(x)}$ denotes the feedback at iteration $x$.
For subsequent iterations ($x\geq 1$), the worker agent refines the output with the input and the feedback.
$Y^{(x)}=A_{W}(C,R^{(x-1)})$ (14)
The feedback taskforce continues until the output meets the manager’s requirements and no further feedback is provided, or until a maximum number of iterations $x_{\max}$ is reached. In the following, we detail how workers cooperate with the manager.
The searching agent formulates search queries based on user instructions, submits queries to search engines, and collects relevant literature. It filters and ranks the retrieved references by citation count, removes incomplete entries, and selects the top $Q_{k}$ unique references for further use. The outlining agent utilizes the gathered literature and user instructions to prompt an LLM for generating sub-outlines. The locating agent downloads and parses full-text papers, applies LLM-based previewing to select pertinent pages, and extracts concise text snippets for drafting. The drafting agent prompts the LLM to synthesize coherent text by incorporating selected references and prior drafts.
@@ 4.1 Experimental Settings
We conduct experiments on two benchmark datasets, AutoSurvey Wang et al. (2024b) and SurveyEval Wang et al. (2025b), as well as a self-constructed benchmark dataset, TopSurvey. For hyperparameters, we set $d_{\max}$, $t_{\max}$, and $x_{\max}$ to 4. The hyperparameter $\theta$ is set to 500. We use GPT-4.1 for all agents. The evaluation comprises two categories of metrics. For citation quality, we adopt citation recall and citation precision as proposed by Wang et al. (2024b): recall measures whether cited passages fully support all statements, while precision measures the proportion of relevant citations that support their corresponding statements. For content quality, following Wang et al. (2024b), we use coverage, structure, and relevance, each rated by LLMs on a 5-point scale. Coverage assesses the extent to which the survey addresses all relevant aspects of the topic; structure evaluates logical organization and coherence; and relevance measures alignment with the specified research topic. We do not filter out fractional scores, such as 4.5.
4.2 Comparision Experiments
@@
We first evaluate our framework on the AutoSurvey Wang et al. (2024b) benchmark, following the protocols established by Liang et al. (2025); Wang et al. (2024b). We use the same 20 topics from diverse subfields of LLM research to generate survey articles for comparison. We compare our approach with three baselines: Naive RAG-based LLMs using retrieval-augmented generation, AutoSurvey Wang et al. (2024b), and SurveyX Liang et al. (2025). We follow the official report of AutoSurvey [^1] https://github.com/AutoSurveys/AutoSurvey and conduct experiments with the official implementation of the offline version’s SurveyX [^2] https://github.com/IAAR-Shanghai/SurveyX.
As shown in Table 1, most algorithms exhibit reduced performance as survey length increases, particularly in recall, precision, structure, and relevance. This decline is most pronounced in the Naive RAG baseline, indicating that longer workflows lead to greater error accumulation and propagation. In contrast, coverage remains stable or improves with longer surveys, likely due to more comprehensive topic inclusion. State-of-the-art methods such as AutoSurvey and SurveyX continue to face challenges with citation precision, frequently generating statements unsupported by references. For content quality, structure consistently receives the lowest scores, reflecting disorganized article organization. Our proposed method remains robust to these issues and consistently achieves superior results across all metrics.
We further evaluate our framework on the SurveyEval benchmark Wang et al. (2025b), using the same protocols. SurveyEval is the first benchmark in computer science that pairs surveys with complete reference papers, comprising 384 arXiv cs.CL surveys citing over 26,000 references. Twenty topics were selected for testing based on reference completeness and reference list diversity. Experimental results are shown in Fig. 3, with Coverage, Structure, and Relevance scores normalized to a 100-point scale.
@@
[…]
Table 4: Sensitivity experiments across different iterations for the experience, exploration, and exploitation taskforces on the new large-scale benchmark.
@@ 4.3 Ablation Study
We further conduct an ablation study to evaluate the effectiveness of each component in our proposed framework. Experiments are performed on the same benchmark dataset as before, with all experimental settings unchanged and the survey length set to 8k tokens. We compare the full model with three variants. We use a single round of retrieval and organization for the variant without the exploration taskforce to generate the overall outline. We draft the manuscript using only one round of extraction and writing for the variant without the exploitation taskforce. For the variant without the feedback taskforce, each agent completes its assigned task, without revision.
Table 2 presents the results. Removing the exploration taskforce causes the largest drops in recall and structure, demonstrating its importance for citation recall and logical organization. Excluding the exploitation taskforce sharply reduces precision and relevance, confirming its role in improving citation precision and content relevance. Without the feedback taskforce, all metrics decrease, especially coverage, highlighting its key role in ensuring comprehensive coverage through collaborative revision.
@@ 4.4 A New Large-Scale Benchmark
We conduct experiments using the same settings as above, generating 64k-token literature reviews for evaluation. As shown in Table 3, compared with results on the existing benchmark in Table 1, performance on the new large-scale benchmark drops significantly, particularly in recall and coverage. This may be due to the greater number and broader range of topics, which leads to some relevant literature not being retrieved, resulting in lower citation recall and coverage. Despite this, our proposed method achieves over 80% in citation scores and an average content quality score of 4.92.
To verify the consistency between LLM evaluation methods and human evaluation, we follow Wang et al. (2024b) to conduct a meta-evaluation. Human experts rank the generated surveys, and we compare these rankings with those generated by LLMs using Spearman’s rank correlation coefficient. The LLM evaluation achieves a correlation of 0.5117, suggesting LLMs align well with human preferences.
4.5 Sensitivity Analysis
To gain deeper insight into our framework, we conduct a sensitivity analysis to observe how $d_{\max}$, $t_{\max}$, and $x_{\max}$ affect performance. We evaluate the experience, exploration, and exploitation taskforces across different numbers of iterations, employing a controlled variable approach: when varying one hyperparameter, all others are held constant as previously described.
The results are presented in Table 4. We find that the second and third iterations yield the most significant improvements. This indicates that errors can accumulate at various steps in the workflow, and iterative refinement effectively reduces intermediate errors, thereby enhancing the quality of the final results. Both the Experience and exploitation taskforces contribute most to faithfulness; from the first to the fourth iteration, the number of references supporting claims in the final output increases by a large margin. Additionally, the Experience and exploration taskforces contribute notably to coverage and structural quality, suggesting that a single revision is often insufficient for comprehensive improvement. However, the effect of additional iterations gradually diminishes. By the fifth iteration, we observe a slight decline, due to unnecessary modifications overwriting correct results and introducing new errors.
Cluster Citation Quality Content Quality Recall $\uparrow$ Precision $\uparrow$ Coverage $\uparrow$ Structure $\uparrow$ Relevance $\uparrow$ 1 83.61 82.57 4.67 4.78 4.93 2 84.54 75.12 4.75 4.85 4.85 3 72.29 76.11 4.59 4.88 4.87 4 86.63 81.98 4.85 4.90 5.00 5 81.37 69.77 4.66 4.74 4.73
Table 5: Real-world case study across different topics on 400 generated results. Higher scores indicate better performance.
4.6 Disscussions
@@
We evaluated our framework in a real-world setting by deploying an online automatic literature review generation system [^3] Anonymous for review; to be revealed upon acceptance., which has produced over 20,000 reviews. We randomly selected 400 reviews, embedded them with all-MiniLM-L6-v2, and applied K-means clustering to form five groups. Results show differences in citation and content quality among clusters. The best performance appears in computer science (cluster 4), while education and social sciences (cluster 3) perform worse, especially in citation quality. It may result from the search engine not including part of the relevant literature in the social sciences. We also measured system efficiency, and generating an 8,000-token review takes an average of 8.45 minutes.
@@ 5 Conclusion
In this paper, we identified and addressed the critical challenge of compounding errors in automated long-form literature review generation. The proposed Multi-Agent Taskforce Collaboration framework strategically orchestrates LLM-based agents through three specialized taskforces. Extensive experiments demonstrate that MATC achieves state-of-the-art performance across established benchmarks. The robustness and generalizability of our framework are further confirmed on our newly contributed TopSurvey benchmark, a large-scale collection of 195 peer-reviewed topics. The deployment of our online system, which has generated over 20,000 reviews, provides practical evidence of MATC’s effectiveness and efficiency in real-world scenarios.

### v3

@@ 3.3 Self-Loop Memory
The Self-Loop Memory is designed to execute tasks assigned by $A_{M}$, taking into account the risk of deviations at each step arising from unpredictable agent-environment interactions.
@@
In detail, for each worker agent $A_{W}$ (which can be $A_{S}$, $A_{O}$, $A_{L}$, or $A_{D}$), $A_{M}$ initiates the Self-Loop Memory in the form of historical experience. $A_{W}$ first generates the primary results, then $A_{M}$ leverages historical experience to review these results, after which $A_{W}$ refines its outputs based on the feedback provided. At the initial step, when the historical experience is empty, $A_{W}$ produces the output solely based on the context:
@@
where $Y^{(0)}$ is the initial output, and $C$ is the context fetched by $A_{M}$.
In the following steps, let $\mathcal{R}$ denote the set of all historical feedback records. For each feedback $r\in\mathcal{R}$, let $\text{revision}(r)$ denote the number of failures and further revisions required after applying feedback $r$, and let $\text{usage}(r)$ denote the number of times feedback $r$ has been selected for demonstration. Let $r_{\min}=\min_{r\in\mathcal{R}}\text{revision}(r)$ denote the minimum number of revisions across all feedback records. We select useful feedback and summarize it into a checklist to guide the worker’s output at each iteration $x$:
$r^{*}=\arg\max_{r\in\mathcal{R}_{\min}}\text{usage}(r)$ (12)
where $\mathcal{R}_{\min}=\{r\in\mathcal{R}:\text{revision}(r)=r_{\min}\}$ is the subset of feedback records that led to the fewest further revisions. Within $\mathcal{R}_{\min}$, we select the most frequently used one, i.e., the one with the highest $\text{usage}(r)$.
Referring to $r^{*}$, $A_{M}$ then provides feedback on the worker’s output:
$R^{(x)}=A_{M}(Y^{(x)},r^{*})$ (13)
where $R^{(x)}$ denotes the feedback at iteration $x$.
For subsequent iterations ($x\geq 1$), $A_{W}$ refines the output with the input and the feedback:
$Y^{(x)}=A_{W}(C,R^{(x-1)})$ (14)
The Self-Loop Memory continues until the output meets the requirements of $A_{M}$ and no further feedback is provided, or until a maximum number of iterations $x_{\max}$ is reached.
In the following, we detail how worker agents cooperate with $A_{M}$. $A_{S}$ formulates search queries based on user instructions, submits queries to search engines, and collects relevant literature. It filters and ranks the retrieved references by citation count, removes incomplete entries, and selects the top $Q_{k}$ unique references for further use. $A_{O}$ utilizes the gathered literature and user instructions to prompt an LLM for generating sub-outlines. $A_{L}$ downloads and parses full-text papers, applies LLM-based previewing to select pertinent pages, and extracts concise text snippets for drafting. $A_{D}$ prompts the LLM to synthesize coherent text by incorporating selected references and prior drafts.
@@ 4.1 Experimental Settings
We conduct experiments on two benchmark datasets, AutoSurvey Wang et al. (2024b) and SurveyEval Wang et al. (2025b), as well as a self-constructed benchmark dataset, TopSurvey. For hyperparameters, we set $d_{\max}$, $t_{\max}$, and $x_{\max}$ to 4. The hyperparameter $\theta$ is set to 500. We use GPT-4.1 Achiam et al. (2023) for all agents. The evaluation comprises two categories of metrics. For citation quality, we adopt citation recall and citation precision as proposed by Wang et al. (2024b). Recall measures whether cited passages fully support all statements. Precision measures the proportion of relevant citations that support their corresponding statements. For content quality, following Wang et al. (2024b); Liang et al. (2025), we use coverage, structure, and relevance, each rated by LLMs on a 5-point scale. Coverage assesses the extent to which the survey encapsulates all relevant aspects of the topic, ensuring comprehensive discussion on both central and peripheral topics. Structure evaluates the logical organization and coherence of sections and subsections, ensuring that they are logically connected. Relevance measures how well the content of the survey aligns with the research topic and maintains a clear focus. We do not filter out fractional scores, such as 4.5. We use GPT-4o Achiam et al. (2023) as the evaluation model for fair comparison Liang et al. (2025).
4.2 Comparison Experiments
@@
We first evaluate our framework on the AutoSurvey Wang et al. (2024b) benchmark, following the protocols established by Liang et al. (2025). We use the same 20 topics from diverse subfields of LLM research to generate survey articles for comparison. We compare our approach with three baselines, including Naive RAG-based LLMs using retrieval-augmented generation, AutoSurvey Wang et al. (2024b), and SurveyX Liang et al. (2025). We follow the official report of AutoSurveyhttps://github.com/AutoSurveys/AutoSurvey and conduct experiments with the official implementation of the offline version of SurveyXhttps://github.com/IAAR-Shanghai/SurveyX.
As shown in Table 1, most methods exhibit reduced performance as survey length increases, particularly in recall, precision, structure, and relevance. This decline is most pronounced in the Naive RAG baseline, indicating that longer workflows lead to greater error accumulation and propagation. In contrast, coverage remains stable or improves with longer surveys, likely due to more comprehensive topic inclusion. State-of-the-art methods such as AutoSurvey and SurveyX continue to face challenges with citation precision, frequently generating statements unsupported by references. For content quality, structure consistently receives the lowest scores, reflecting disorganized article organization. Our proposed method remains robust to these issues and consistently achieves superior results across all metrics.
We also compare with recent state-of-the-art LLMs, including DeepSeek-R1 Guo et al. (2025), Gemini-3 Team et al. (2023), and GPT-5.2. These recent LLMs possess more up-to-date knowledge and enhanced reasoning capabilities. We equip them with an academic search tool, enabling them to retrieve relevant literature when necessary for generating the review. Even with paid subscriptions, state-of-the-art LLMs do not consistently generate long-form content exceeding 8k tokens as instructed, consistent with recent observations on the length control challenge Akinfaderin et al. . Therefore, we report results only at the 8k-token length. The results show that Gemini-3 and GPT-5.2 demonstrate comparable performance to the agent system with GPT-4.1, highlighting the importance of the underlying foundation model.
We further evaluate our framework on the SurveyEval benchmark Wang et al. (2025b), using the same protocols. SurveyEval is the first benchmark in computer science that pairs surveys with complete reference papers, comprising 384 arXiv cs.CL surveys citing over 26,000 references. Twenty topics are selected for testing based on reference completeness and reference list diversity. Experimental results are shown in Fig. 3, with coverage, structure, and relevance scores normalized to a 100-point scale.
@@
Cluster Citation Quality Content Quality Recall $\uparrow$ Precision $\uparrow$ Coverage $\uparrow$ Structure $\uparrow$ Relevance $\uparrow$ Avg. $\uparrow$ Physics 82.94 81.43 4.69 4.76 4.91 4.79 Basic Medicine 83.87 76.35 4.73 4.83 4.84 4.80 Education 73.52 75.68 4.61 4.85 4.84 4.77 Computer Science & Technology 85.78 81.26 4.82 4.88 4.97 4.89 Economics 80.91 71.24 4.68 4.75 4.76 4.73
Table 4: Real-world case study across different topics on 10,000 generated reviews. Higher scores indicate better performance.
@@ 4.3 Ablation Study
We conduct an ablation study to evaluate the effectiveness of each component in our proposed framework. Experiments are performed on the same benchmark dataset as before, with all experimental settings unchanged and the survey length set to 8k tokens. We compare the full model with three variants. For the variant without Tree-Guided Memory, we use a single round of retrieval and organization to generate the overall outline. For the variant without Hub-Guided Memory, we draft the manuscript using only one round of extraction and writing. For the variant without Self-Loop Memory, each agent completes its assigned task without revision.
Table 2 presents the results. Removing Tree-Guided Memory causes the largest drops in recall and structure, demonstrating its importance for citation recall and logical organization. Excluding Hub-Guided Memory sharply reduces precision and relevance, confirming its role in improving citation precision and content relevance. Without Self-Loop Memory, all metrics decrease, especially coverage, highlighting its key role in ensuring comprehensive coverage through collaborative revision.
@@ 4.4 A New Large-Scale Benchmark
We conduct experiments using the same settings as above, generating 64k-token literature reviews for evaluation. As shown in Table 3, compared with the results on the existing benchmark in Table 1, performance on the new large-scale benchmark drops significantly, particularly in recall and coverage. This may be due to the greater number and broader range of topics, which leads to some relevant literature not being retrieved, resulting in lower citation recall and coverage. Despite this, our proposed method achieves over 80% in citation scores and an average content quality score of 4.92.
@@
Finally, we evaluate our framework in a real-world setting by deploying an online automatic literature review generation systemAnonymous for review; to be revealed upon acceptance., which has produced approximately 40,000 reviews to date. We randomly sampled 10,000 reviews, embedded them using all-MiniLM-L6-v2, and clustered them into five groups corresponding to the five most frequently requested disciplines. As shown in Table 4, results reveal notable differences in citation and content quality across clusters. The best performance is observed in Computer Science & Technology, while Education performs comparatively worse, particularly in citation quality. This discrepancy may be attributed to the search engine’s limited coverage of relevant literature in the social sciences. We also measured system efficiency: generating an 8,000-token review takes an average of 8.45 minutes. To facilitate future research, we release the titles and instructions of the 10,000 sampled reviews in English for comparative studies.
5 Discussion
Our experimental results demonstrate that SGMC outperforms not only pipeline-based agents but also the latest standalone LLMs equipped with retrieval tools. This raises a crucial question: in the era of increasingly powerful foundation models, what is the unique value of a structured framework like SGMC? Our structured memory mechanism offers two key advantages over pure LLM-based approaches. First, the tree-guided outline and hub-guided memory nodes create an explicit, traceable chain from any final claim back to its source literature, enabling errors to be pinpointed to specific retrieval, outlining, or drafting steps. Standalone LLMs lack this transparent provenance, making error diagnosis and systematic correction unreliable. Second, SGMC embeds proactive verification within the workflow through Self-Loop Memory, preventing errors from propagating to subsequent stages. In contrast, methods applicable to standalone LLMs, such as self-refinement or post-editing, are inherently reactive and often insufficient for correcting deeply embedded, compounded errors. The ablation study further substantiates that each component of our structured memory is vital for holistic quality. We acknowledge the remarkable capabilities of state-of-the-art LLMs, and the agents within SGMC can be instantiated by these powerful models. The contribution of SGMC is to provide a reliable, controllable, and error-aware process framework that channels the generative power of LLMs toward producing high-fidelity, verifiable scholarly content. For automated literature review generation, this structural guidance is essential.
@@ 6 Conclusion
In this paper, we tackle the challenge of compounding errors in automated long-form literature review generation. We propose Structure-Guided Memory Consolidation (SGMC), a framework that systematically consolidates and verifies information throughout the multi-stage generation process using tree-guided, hub-guided, and self-loop memory modules. Extensive experiments demonstrate that SGMC achieves state-of-the-art performance in citation accuracy and content quality. Moreover, the successful deployment of our online system, which has generated over 40,000 reviews, provides strong evidence of SGMC’s practical effectiveness.